Grade 11Mechanics

momentum of a particle is increased by 50%.By how much percentagekinetic energy of particle will increase?

RAJAT  PORWAL
14 Years agoGrade 11

2 Answers

Ashwin Muralidharan IIT Madras
ApprovedApproved Tutor Answer14 Years ago

Hi Rajat,

 

As Momentum = MxV, means when momentum is increased by 50%, velocity is increased by 50%

So if v1 was the intial velocity. Now v2 = 1.5V is the final Velocity.

 

Initial KE = (1/2m)v12 = k1

Final KE = (1/2m)v22 = k2

 

k2/k1 = v22/v12 = 1.52 = 2.25

 

Hence the percentage increase in KE = (2.25-1)*100% = 125%

 

Hope that helps.

 

All the best,

Regards,

Ashwin (IIT Madras).

suchita undare
14 Years ago

let the momemtum be p

inc by 50% i.e p+p/2=3p/2

K.Ei = P2/2m    K.Ef = (3P/2)2/2m=9P2/8m

change in K.E = 5P2/8m

% inc =(change in K.E/initial K.E)multiplied by 100=75%